OpenStudy (anonymous):

parking in a student lot costs $1 for the first half hour, and 1.75 for each hour there after. A partial hour is charged the same as a full hour. What is the longest time that a student can park in the lot for $8.00

OpenStudy (anonymous):

Make a list of how many hours you park, and how much it will cost.

OpenStudy (anonymous):

grind it til you find it up to one half hour $1 up to one and a half hours $2.75 up to two and a half hours $4.50 etc

OpenStudy (anonymous):

Rule: Charged by full hours, no partial hours. Limit of $8. Convert Rates to a common unit, Hours: 1.) First 30 minutes (x): $1*2 = $2/hr 2.) Remaining time (y): $1.75/hr We will find that t= x+y x= 30min./60min. = .5 y=unknown $8 = The most we can pay. So we know this: 1.) $8 = $2*(x) + $1.75*(y), where x=.5hrs and y=remaining time 2.) $8 = $2*(.5) + $1.75*(y) 3.) $8 = $1 + $1.75*(y), now we subtract $1 from both sides -$1 -$1 4.) $8-$1 = $1.75*(y), we get this equation afterward 5.) $7 = $175(y), now we divide by $1.75 on both sides /$1.75 /$1.75 6.) $7/$1.75 = (y), we get this equation afterward 7.) $7/$1.75 = 4, so y=4 So now we know x=.5 and y=4. Our total time equation is t=x+y. So, let's plug in our values for x and y. 1.) t = x+y 2.) t = .5 + 4 3.) 4.5 = .5 + 4 But we have to pay the full hour because there are no partial hours. So, 4.5 would cost the same as 5 hours. Since we don't have anymore money, we can't pay for 5 hours of parking. We need to round down to the nearest full hour. In this case, rounding down will give us 4 hours. So our answer is 4 hours of parking. And, we will have $1.75*(.5)=$87.5 cents left in our wallet. I hope that helps.

OpenStudy (anonymous):

make a proportion